If you dilute 20.0 L of a 4.5 M solution to make 150 L of solution, what is the molarity of the dilute solution?
1
Expert's answer
2018-03-29T07:56:44-0400
The amount of moles in the initial and resulting solutions is constant.
V1 ·CM1 = V2·CM2
20 · 4.5 = 150 · CM2
CM2 = (20·4.5)/150 = 0.6 M